# Mind: Towards Immersive Psychological Healing with Multi-agent Inner Dialogue

### v1

@@ Abstract
Mental health issues are worsening in today’s competitive society, such as depression and anxiety. Traditional healings like counseling and chatbots fail to engage effectively, they often provide generic responses lacking emotional depth. Although large language models (LLMs) have the potential to create more human-like interactions, they still struggle to capture subtle emotions. This requires LLMs to be equipped with human-like adaptability and warmth.
To fill this gap, we propose the Mind (Multi-agent INner Dialogue), a novel paradigm that provides more immersive psychological healing environments. Considering the strong generative and role-playing ability of LLM agents, we predefine an interactive healing framework and assign LLM agents different roles within the framework to engage in interactive inner dialogues with users, thereby providing an immersive healing experience. We conduct extensive human experiments in various real-world healing dimensions, and find that Mind provides a more user-friendly experience than traditional paradigms. This demonstrates that Mind effectively leverages the significant potential of LLMs in psychological healing.
@@
Yujia Chen1, Changsong Li1†, Yiming Wang2, Qingqing Xiao3,4, Nan Zhang1, Zifan Kong1, Peng Wang1, Binyu Yan1††thanks: Corresponding author. 1Sichuan University 2Shanghai Jiao Tong University 3Mental Health Center, West China Hospital, Sichuan University 4WestChina School of Nursing, Sichuan University {yujiachen0058,changsongli0018,alsaceym}@gmail.com {yanby}@scu.edu.cn
@@ 1 Introduction
Mental health issues are worsening in today’s competitive society, with rising cases of disorders like depression (Moitra et al., 2023). This lead to a growing market for psychological healing. Traditional healing paradigms like Cognitive Behavioral Therapy (Beck, 1979) and Dialectical Behavior Therapy (Lynch et al., 2007) are widely used but rely on face-to-face interactions, making them time-consuming and costly (Duruz et al., 2003) that limits large-scale accessibility. Another healing paradigm is VR-based Empathy Training (Halim et al., 2023; Hidding et al., 2024; Döllinger et al., 2024), which is rooted in the mechanisms of virtual body ownership illusion. This method involves the creation of a virtual self-image within VR, where the individual provides verbal encouragement to this virtual self. Subsequently, the individual swaps perspectives to adopt the role of the one being comforted, receiving the previously offered words of comfort. This “self-dialogue” process enables patients to experience and understand empathy from a new vantage point, thereby promoting self-compassion and reducing self-criticism. However, current systems are limited by static scenarios and scripted interactions. The lack of counselor-guided mechanisms and dynamic responsiveness in these predefined frameworks hampers the effectiveness of emotional regulation and reduces the adaptability of the therapeutic process, potentially leading to emotional deterioration in self-administered interventions.
@@
Based on the above motivations, we propose Multi-agent INner Dialogue (Mind), a novel immersive and interactive psychological healing paradigm. As illustrated in Figure 1, our approach is the first to introduce a multi-agent system into an empathy training paradigm, significantly enhancing the interaction between the patient and their inner self through dynamic narrative scenarios. We first decompose the patient’s conflicting self into interactive conversational entities, creating an embodied space for inner dialogue, and introducing a dynamic scene generation mechanism that adapts the narrative path based on the user’s emotional state. Next, we allocate four core roles for LLM agents, including trigger, devil, guide, and strategist. Through multi-perspective dialogue history review, it provides the user with a cognitive scaffold for metacognitive reflection. This design allows the patient to not only experience empathy transformation from a dual perspective but also, through the agent-mediated dialogue process, systematically deconstruct the cognitive roots of emotional generation, leading to deeper self-reconciliation.
We conduct extensive experiments to validate our approach. Initially, we benchmark the performance of seven large language models (LLMs) in simulating patient roles. The main experiment involved a comparative evaluation where human evaluators assessed Mind against traditional psychologist counseling, chatbots, and conventional empathy training methods. The results demonstrate that Mind has the potential to advance psychological interventions by combining the scalability of LLMs with human-centered interaction design, thereby offering accessible and effective mental health support. Additional experiments based on different themes and ablation experiments based on agents further underscored the stability in generated scenarios and content and the rationality of the overall framework.
@@ 2 Mind: Multi-agent Inner Dialogue
2.1 Overall Framework
The overall framework of our Mind paradigm is shown in Figure 2, composed of four agents responsible for inner dialogue generation, in addition to an agent simulating patients with cognitive distortions. The subsequent section will commence with an overview of the workflow: the trigger, the devil, the guide, the strategist and the human simulated patient. Detailed prompt templates used by each agent are presented in Appendix C.
In the initial phase of the psychological interaction framework, the player articulates their recent concerns and selects a theme to guide the narrative direction. These inputs serve as the foundational trigger for dynamically generating virtual scenarios (), which are designed to mirror the player’s cognitive patterns. Subsequently, the devil processes the scenario () and concerns to distorted thoughts (), emulating maladaptive cognitive biases aligned with the player’s mental state.
The guide then integrates and to generate professional psychological guidance (), aimed at facilitating empathetic responses from the player. Upon receiving , the player engages in a reflective dialogue to provide comforting words () and counter the distorted thoughts (), thereby advancing the therapeutic narrative.
To ensure narrative coherence and progression, the strategist analyzes the cumulative memory () — a structured summary of prior scenarios, cognitive distortions, and guidance — alongside the player’s comforting words (). This analysis produces strategic directives () that govern the generation of subsequent triggers () and the devil’s adaptive cognitive evolution ().
Through iterative cycles of scenario generation, cognitive reflection, and guided intervention, the framework progressively refines its alignment with the player’s psychological profile. Each iteration enhances the system’s capacity to model nuanced mental states while maintaining narrative continuity, ultimately achieving a balance between therapeutic efficacy and immersive storytelling.
@@ 2.2 Trigger: Scenario Generation
where is the player’s concerns and is the theme, is the player’s last-round comforting words and is the strategist’s last-round storyline progression.
@@ 2.3 Devil: Cognitive Distortion Simulation
The devil simulates the cognitive distortions that a patient might experience within the context of the scenario. It functions as the player’s “virtual embodiment” representing an “alternate self” within the simulated environment.
@@ 2.4 Guide: Cognitive Restructuring Guidance
As the game progresses, the growing history becomes burdensome for the LLM to process efficiently. To mitigate this issue, a summarization mechanism is employed to maintain coherent narrative memory (Zhou et al., 2023). By operating in this way, the guide ensures that the player is not only challenged but also supported in a structured, manageable way, encouraging long-term emotional resilience and rational thinking. Ultimately, the guide helps transform the player from a passive recipient of distorted thoughts, as influenced by the devil, into an active participant in their own cognitive change, laying the foundation for healthier thought patterns and emotional well-being.
@@ 2.5 Strategist: Storyline Progression
where is the player’s this round comforting words.
@@ 2.6 Human Simulated Patient: Empathy and Interaction
To facilitate the automated operation and evaluation of our framework, and drawing upon the validated psychological characteristics and annotation capabilities of LLM, we employ LLMs to simulate human interactions by providing comforting words to the devil. Based on the guidance from the guide, the virtual scenario generated by the trigger, and the cognitive distortions produced by the devil, human simulated patient assumes the role of the Player, engaging in empathetic reassurance toward the devil . This process also incorporates the chain-of-thought (CoT) technique(Wei et al., 2022), allowing for a structured and coherent response generation that aligns with the psychological progression of the player-agent interaction.
@@ 3 Experiments
| Metric | Description |
| Immersion | Measures whether the user feels fully engaged and captivated by the interaction. |
| Coherence | Assesses if the generated content is logical and transitions smoothly. |
| Engagement | Evaluates if the system encourages sustained and meaningful interaction. |
| Emotional Relief | Measures if the interaction reduces user stress or anxiety. |
| Satisfaction | Reflects the user’s overall contentment with the system. |
| Interest | Assesses whether the content grabs attention and sparks curiosity. |
@@ 3.1 Setup
The real-life scenarios, thinking patterns, and cognitive distortion types of the Human Simulated Patient simulated by the LLM are derived from the C2D2 dataset (Wang et al., 2023). This dataset is the first publicly available resource focused on cognitive distortion analysis, solving the problem of data scarcity in this field. The dataset covers eight major topics, including work issues, interpersonal issues, economic issues, random negative events, family issues, physical stress, and discrepancy between ideal and reality.
@@
We used several LLM agents including both open-source and closed-sourced models with varying parameter scales. For closed-source models, we chose Gemini-2.0-flash (gem, 2025), GPT-4o (OpenAI et al., 2024), GPT-3.5-Turbo (Ye et al., 2023). For open-source models, we chose Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Qwen2.5-72B-Instruct (Qwen et al., 2025), Qwen2.5-7B-Instruct (Qwen et al., 2025) and Deepseek-R1 (DeepSeek-AI et al., 2025). To control variables, we set the temperature of each model to 0.7.
In Section 3.2, a preliminary role-playing experiment was conducted to evaluate the performance of various models in the SP role-playing task. Based on the results of this initial assessment, we selected the Gemini-2.0-flash model, which exhibited the best performance, for our main experiments. This model was chosen due to its superior ability to handle the complexities of the role-playing scenario, making it the most appropriate candidate for further investigation in the primary phase of our study.
@@
The generation quality of the devil agent is crucial for the implementation of this framework, as it mirrors the player’s internal “cognitive distortions” and can better serve its purpose only if it closely aligns with the player’s “inner voice.” Therefore, before conducting our main experiments, we have specifically designed an Simulated Patient(SP) role-playing evaluation system to verify that the model can correctly identify the player’s type of cognitive distortion and accurately reflect the patient’s thoughts, achieving a realistic effect. We recruited 5 mental health professionals. Each evaluator had 10 rounds of conversation with each model. The evaluators rated the content based on the five evaluation metrics (Johri et al., 2025), with a scoring range of 1 to 5. Detailed illustration is shown in Appendix B.
The paradigm comparison evaluation metrics we propose cover three main aspects: user experience, interaction quality, and emotional comfort, with six different metrics (Hua et al., 2024; Kumaran et al., 2023; Jennett et al., 2008; Ryan, 2015; Nacke and Drachen, 2011). The evaluation metrics is shown in Table 1. We recruited 7 mental health practitioners with professional expertise in psychological therapy. For the different paradigms, the evaluators rated the content based on the six evaluation metrics, with a scoring range of 1 to 5.
| Model Name | DS | LF | EE | PD | Acc |

### v2

@@ Abstract
Mental health issues are worsening in today’s competitive society, such as depression and anxiety. Traditional healings like counseling and chatbots fail to engage effectively, they often provide generic responses lacking emotional depth. Although large language models (LLMs) have the potential to create more human-like interactions, they still struggle to capture subtle emotions. This requires LLMs to be equipped with human-like adaptability and warmth. To fill this gap, we propose the Mind (Multi-agent INner Dialogue), a novel paradigm that provides more immersive psychological healing environments. Considering the strong generative and role-playing ability of LLM agents, we predefine an interactive healing framework and assign LLM agents different roles within the framework to engage in interactive inner dialogues with users, thereby providing an immersive healing experience. We conduct extensive human experiments in various real-world healing dimensions, and find that Mind provides a more user-friendly experience than traditional paradigms. This demonstrates that Mind effectively leverages the significant potential of LLMs in psychological healing.
@@
Yujia Chen1111The two authors contribute equally., Changsong Li1111The two authors contribute equally., Yiming Wang2, Tianjie Ju2,5, Qingqing Xiao3,4, Nan Zhang1, Zifan Kong1, Peng Wang1, Binyu Yan1222Corresponding author. 1Sichuan University 2Shanghai Jiao Tong University 3Mental Health Center, West China Hospital, Sichuan University 4WestChina School of Nursing, Sichuan University 5National University of Singapore {yujiachen0058,changsongli0018,alsaceym}@gmail.com {yanby}@scu.edu.cn
@@ 1 Introduction
Mental health issues are worsening in today’s competitive society, with rising cases of disorders like depression (Moitra et al., 2023). This lead to a growing market for psychological healing. Traditional healing paradigms like Cognitive Behavioral Therapy (Beck, 1979) and Dialectical Behavior Therapy (Lynch et al., 2007) are widely used but rely on face-to-face interactions, making them time-consuming and costly (Duruz et al., 2003) that limits large-scale accessibility.
Another healing paradigm, VR-based Empathy Training (Halim et al., 2023; Hidding et al., 2024; Döllinger et al., 2024), involves self-dialogue in virtual reality, where individuals alternate perspectives between comforting and being comforted through a virtual self-representation. This process enhances self-empathy, thereby promoting self-compassion and reducing self-criticism. However, current systems are limited by static scenarios and scripted interactions. The absence of counselor guidance and flexible feedback in these fixed frameworks limits emotional regulation and weakens the adaptability of therapy.
@@
Based on the above motivations, we propose Multi-agent INner Dialogue (Mind), a novel immersive and interactive psychological healing paradigm. As illustrated in Figure 1, our approach is the first to introduce a multi-agent system into an empathy training paradigm, significantly enhancing the interaction between the user and their inner self through dynamic narrative scenarios. Unlike traditional approaches that attempt to directly establish empathic alliances between therapists (or LLMs) and users, our framework emphasizes indirectly cultivating self-compassion in self-critical individuals. By assigning empathetic roles to LLM agents and directing empathy towards the user’s own “inner self,” our design enables users to engage in a simulated yet emotionally resonant “self-to-self” caring process. This mechanism is particularly beneficial for individuals who struggle to generate self-compassion through imagination or guided visualization. We allocate four core roles to LLM agents (Trigger, Devil, Guide, and Strategist) each serving distinct reflective and emotional functions. This design helps foster internal empathy shifts and reveals cognitive patterns behind distress, promoting self-acceptance and psychological healing.
We conduct extensive experiments, including human evaluations, human experiments and ablation studies. The results demonstrate that:
Mind outperforms traditional counseling, chatbots, and traditional empathy training methods, achieving an average improvement of 13% across six psychological dimensions.
Mind demonstrates the highest positive emotional shift (1.46) and largest negative affect reduction (-0.65) among all dialogue systems, while also receiving the highest overall ratings from participants.
Ablation studies emphasize the significance of the memory mechanism, guide agent, and strategist agent, with an average performance drop of 42% when these components are removed.
@@ 2 Mind: Multi-agent Inner Dialogue
2.1 Overall Workflow
The overall framework of our Mind paradigm is shown in Figure 2, composed of four agents responsible for inner dialogue generation, in addition to an agent simulating patients with cognitive distortions. The subsequent section will commence with an overview of the workflow: the trigger, the devil, the guide, the strategist and the human simulated patient. Detailed prompt templates used by each agent are presented in Appendix F.
In this framework, stands for the -th interaction. denotes the virtual scenarios. represents the distorted thoughts. refers to the professional psychological guidance. indicates the comforting words provided by the player. is the cumulative memory, which is a structured summary of previous scenarios, cognitive distortions. represents the storyline progression. Additionally, stands for the player’s concerns, and denotes the overarching theme of the interaction.
Step 0: In the initial turn, the player articulates their current concern and selects a theme , which together serve to anchor and guide the direction of the narrative.
Step 1: The trigger is dynamically generated based on either the player’s initial inputs and (in the first turn), or the reflective response and planning signal (in subsequent turns). is constructed to reflect the player’s concerning scenes.
Step 2: The devil processes , and to , emulating maladaptive cognitive biases aligned with the player’s mental state.
Step 3: The guide then integrates and to generate , aimed at facilitating empathetic responses from the player.
Step 4: Upon receiving , the player engages in a reflective dialogue to provide and counter , thereby advancing the therapeutic narrative.
Step 5: The strategist analyzes and . This analysis produces that govern the generation of subsequent triggers () and the devil’s adaptive cognitive evolution ().
Through iterative cycles of scenario generation, cognitive reflection, and guided intervention, the framework progressively refines its alignment with the player’s psychological profile. The entire algorithm is provided in the Algorithm 1.
@@ 2.3 Devil: Cognitive Distortion Simulation
The devil simulates the cognitive distortions that a patient might experience within the context of the scenario and it is aligned with the concept Simulated Patient(SP). It functions as the player’s “virtual embodiment” representing an “alternate self” within the simulated environment.
@@ 2.4 Guide: Cognitive Restructuring Guidance
As the game progresses, the growing history becomes burdensome for the LLM to process efficiently. To mitigate this issue, a summarization mechanism is employed to maintain coherent narrative memory (Zhou et al., 2023). In our implementation, we use a well-designed prompt template (see Appendix F for details) to guide the model to extract key events, emotional states, and cognitive distortion patterns from the history of the interaction and save them to the memory unit. The system recursively compares new and old memories, merges redundant information, and retains core therapeutic cues such as “from self-denial to initial reflection” to ensure coherent and streamlined historical memory. By utilizing this summarization mechanism, the guide ensures that the player is not only challenged but also supported in a structured, manageable way, encouraging long-term emotional resilience and rational thinking. Ultimately, the guide helps transform the player from a passive recipient of distorted thoughts, as influenced by the devil, into an active participant in their own cognitive change, laying the foundation for healthier thought patterns and emotional well-being.
@@ 2.6 Human Simulated Patient: Empathy and Interaction
To facilitate the automated operation and evaluation of our framework, and drawing upon the validated psychological characteristics and annotation capabilities of LLM, we employ LLMs to simulate players with cognitively distorted thinking and provide comforting words to the devil. Based on the guidance from the guide, the virtual scenario generated by the trigger, and the cognitive distortions produced by the devil, human simulated patient assumes the role of the Player, engaging in empathetic reassurance toward the devil. This process also incorporates the chain-of-thought (CoT) technique(Wei et al., 2022), allowing for a structured and coherent response generation that aligns with the psychological progression of the player-agent interaction.
@@ 3.1 Setup
The real-life scenarios, thinking patterns, and cognitive distortion types of the Human Simulated Patient simulated by the LLM are derived from the C2D2 dataset (Wang et al., 2023). This dataset is the first publicly available resource focused on cognitive distortion analysis, solving the problem of data scarcity in this field. The dataset covers eight major topics, including work issues, interpersonal issues, economic issues, random negative events, family issues, physical stress, and discrepancy between ideal and reality. All the experimental results in the body part were conducted in English, and the Chinese experimental results are presented in Appendix D.
@@
We used several LLM agents including both open-source and closed-sourced models with varying parameter scales. For closed-source models, we chose Gemini-2.0-flash (gem, 2025), GPT-4o (OpenAI et al., 2024), GPT-3.5-Turbo (Ye et al., 2023). For open-source models, we chose Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Qwen2.5-72B-Instruct (Qwen et al., 2025), Qwen2.5-7B-Instruct (Qwen et al., 2025) and Deepseek-R1 (DeepSeek-AI et al., 2025). We set the temperature of each model to 0.7.
@@
The quality of the devil agent’s responses is critical to this framework, as it reflects the player’s internal “cognitive distortions” and must closely align with their “inner voice”. To ensure this, we first perform a preliminary SP role-playing evaluation in Section 3.2 to assess whether the model can accurately identify the type of cognitive distortion of the player and realistically express their thoughts. We invited five mental health professionals, each of whom held 10 dialogue rounds with each model. They rated the responses using five evaluation metrics (Johri et al., 2025), on a scale from 1 to 5. Detailed evaluation metrics are shown in Appendix B. Based on this experiment, we will select the best-performing model to conduct our main experiments.
In Section 3.3, we conduct our main experiments to compare our Mind with other paradigms. We evaluate three main aspects: user experience, interaction quality, and emotional comfort, with six different metrics (Hua et al., 2024; Kumaran et al., 2023; Jennett et al., 2008; Ryan, 2015; Nacke and Drachen, 2011). Metric details are shown in Appendix E. We recruited 7 mental health professionals with professional expertise in psychological therapy. For the different paradigms, the evaluators rated the content based on the six evaluation metrics, with a scoring range of 1 to 5.
Model Name | DS | CF | EE | PD | Acc |